I need help solving this problem. It is 9-6=8-1(9+a)
Answer by checkley71(8403)

9-6=8-1(9-A)
3=8-9+A
3=-1+A
A=3+1
A=4 ANSWER.
PROOF:
9-6=8-(9-4)
3=8-5
3=3